# Yoga for Fourier–Mukai partnership

### v2

@@ 4. Fully faithfulness & equivalences
$\mathbf{R}(t_{2})_{\ast}$ $\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(E)$
$=\mathbf{R}(t_{2})_{\ast}\mathbf{R}(g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
$\cong\mathbf{R}(t_{2}\circ g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
$\cong\mathbf{R}(f_{1}^{\prime}\circ t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$
$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)\bigg)$ ([GW23, Proposition 22.84])
$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{R}(t^{\prime})_{\ast}\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K\bigg)$
$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\bigg(\mathbf{L}(f^{\prime}_{2})^{\ast}\mathbf{R}(t_{1})_{\ast}E\otimes^{\mathbf{L}}K\bigg)$
$=\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}(E).$
@@
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. If $\Phi_{K}$ is fully faithful on $D_{\operatorname{qc}}$, then $\mathbf{L}t^{\ast}_{1}$ preserves the unit $\eta$ of the adjoint pair $\Phi_{K}$ and $\Phi_{K^{\prime}}$ on $D_{\operatorname{qc}}$. That is, $\mathbf{L}t_{1}^{\ast}\eta_{E}$ is the unit of $\mathbf{L}t_{1}^{\ast}E$ for the adjoint pair $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ and $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.

### v3

@@ 4. Fully faithfulness & equivalences
$\Phi_{K}\circ\mathbf{R}(t_{1})_{\ast}(E)$ $=\mathbf{R}(f_{1}^{\prime})_{\ast}\big(\mathbf{L}(f^{\prime}_{2})^{\ast}\mathbf{R}(t_{1})_{\ast}E\otimes^{\mathbf{L}}K\big)$
$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\big(\mathbf{R}t^{\prime}_{\ast}\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}K\big)$ (flat base change)
$\cong\mathbf{R}(f_{1}^{\prime})_{\ast}\mathbf{R}t^{\prime}_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$ (projection formula)
$\cong\mathbf{R}(t_{2})_{\ast}\mathbf{R}(g^{\prime}_{1})_{\ast}\big(\mathbf{L}(g^{\prime}_{2})^{\ast}E\otimes^{\mathbf{L}}\mathbf{L}(t^{\prime})^{\ast}K\big)$ (pseudofunctoriality of $\mathbf{R}(-)_{\ast}$)
$=\mathbf{R}(t_{2})_{\ast}\circ\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(E).$
Lemma 4.5.
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then we have a morphism of adjoint pairs as depicted in the diagram
$D_{\operatorname{qc}}(Y_{1})$$D_{\operatorname{qc}}(Y_{2})$$D_{\operatorname{qc}}(Y_{1}\times_{S}T)$$D_{\operatorname{qc}}(Y_{2}\times_{S}T)$$\scriptstyle{\Phi_{K}}$$\scriptstyle{\mathbf{L}t_{1}^{\ast}}$$\scriptstyle{\mathbf{L}t_{2}^{\ast}}$$\scriptstyle{\mathbf{R}(t_{1})_{\ast}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}}$$\scriptstyle{\mathbf{R}(t_{2})_{\ast}}\scriptstyle{\dashv}\scriptstyle{\dashv}$
with left and right comparison transformations given respectively by the isomorphisms $\beta^{K}$, $\alpha^{K}$ of Lemmas˜4.4 and 3.4.
Proof.
Denote by $\xi^{i}\colon\mathbf{L}t_{i}^{\ast}\mathbf{R}(t_{i})_{\ast}\to\operatorname{id}$ for the counit of the adjunction $\mathbf{L}t_{i}^{\ast}\dashv\mathbf{R}(t_{i})_{\ast}$ with $i=1,2$. To prove the desired claim, it suffices to verify that Item˜4.1(1) holds; that is, that the following diagram commutes:
$\mathbf{L}t_{2}^{\ast}\Phi_{K}\mathbf{R}(t_{1})_{\ast}$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\mathbf{L}t_{1}^{\ast}\mathbf{R}(t_{1})_{\ast}$$\mathbf{L}t_{2}^{\ast}\mathbf{R}(t_{2})_{\ast}\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$$\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}.$$\scriptstyle{\beta^{K}_{\mathbf{R}(t_{1})_{\ast}}}$$\scriptstyle{\mathbf{L}t_{2}^{\ast}(\alpha^{K})}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}(\xi^{1})}$$\scriptstyle{\xi^{2}_{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}}}$
For brevity sake, in the remainder of the proof, we drop the $\mathbf{L}$’s and $\mathbf{R}$’s in the notation for the derived functors (i.e. all functors now are understood to be derived). Hence, the diagram above is now
(4.2) $t_{2}^{\ast}(t_{2})_{\ast}\Phi_{(t^{\prime})^{\ast}K}$$t_{2}^{\ast}\Phi_{K}(t_{1})_{\ast}$$\Phi_{(t^{\prime})^{\ast}K}.$$\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}(t_{1})^{\ast}$$\scriptstyle{\xi^{2}_{\Phi_{(t^{\prime})^{\ast}K}}}$$\scriptstyle{t_{2}^{\ast}(\alpha^{K})}$$\scriptstyle{\beta^{K}_{(t_{1})_{\ast}}}$$\scriptstyle{\Phi_{(t^{\prime})^{\ast}K}(\xi^{1})}$
Choose $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$. If we expand the definition of $\alpha^{K}$ and $\beta^{K}$ in Diagram˜4.2 (see the proofs of Lemmas Lemmas˜4.4 and 3.4), we obtain a large diagram consisting of various faces. In what follows, we make this a bit more explicit. Particularly, we spell out each face needed to understand Diagram˜4.2. To describe said diagram more explicitly, we use the following abbreviations for the (natural) isomorphisms:
  • $\operatorname{BC}_{i}$, $i\in\{1,2\}$ for the flat base change isomorphisms
  • $\operatorname{PF}$ for the pojection formula
  • $F_{(-)_{\ast}}$ for functoriality of $(-)_{\ast}$
  • $F_{(-)^{\ast}}$ for functoriality of $(-)^{\ast}$
  • $M_{(-)^{\ast}}$ for monoidality of $(-)^{\ast}$
  • $\xi^{\prime}$ for the counit of $(t^{\prime})^{\ast}\dashv(t^{\prime})_{\ast}$.
Now, consider the following diagrams:
(4.3) $t_{2}^{\ast}(f_{1}^{\prime})_{\ast}((f_{2}^{\prime})_{\ast}(t_{1})_{\ast}E\otimes K)$$t_{2}^{\ast}(f_{1}^{\prime})_{\ast}((t^{\prime})_{\ast}(g_{2}^{\prime})^{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}((f_{2}^{\prime})^{\ast}(t_{1})_{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}(t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes K)$$\scriptstyle{\operatorname{BC}_{1}}$$\scriptstyle{\operatorname{BC}_{2}}$$\scriptstyle{\operatorname{BC}_{2}}$$\scriptstyle{\operatorname{BC}_{1}}$
(4.4) $t_{2}^{\ast}(f_{1}^{\prime})_{\ast}((t^{\prime})_{\ast}(g_{2}^{\prime})^{\ast}E\otimes K)$$t_{2}^{\ast}(f_{1}^{\prime})_{\ast}(t^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}(t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}t_{\ast}^{\prime}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$\scriptstyle{\operatorname{PF}}$$\scriptstyle{\operatorname{BC}_{2}}$$\scriptstyle{\operatorname{BC}_{2}}$$\scriptstyle{\operatorname{PF}}$
(4.5) $(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}((f_{2}^{\prime})^{\ast}(t_{1})_{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}(t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}((t^{\prime})^{\ast}(f_{2}^{\prime})^{\ast}(t_{1})_{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}((t^{\prime})^{\ast}t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$\scriptstyle{\operatorname{BC}_{1}}$$\scriptstyle{M_{(-)^{\ast}}}$$\scriptstyle{M_{(-)}^{\ast}}$$\scriptstyle{\operatorname{BC}_{1}}$
(4.6) $(g_{1}^{\prime})_{\ast}((t^{\prime})^{\ast}t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}(t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}t_{\ast}^{\prime}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$\scriptstyle{\xi^{\prime}}$$\scriptstyle{M_{(-)}^{\ast}}$$\scriptstyle{\operatorname{PF}}$$\scriptstyle{\xi^{\prime}}$
(4.7) $t_{2}^{\ast}(f_{1}^{\prime})_{\ast}(t^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$t_{2}^{\ast}(t_{2})_{\ast}(g_{1}^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}(t^{\prime})^{\ast}t_{\ast}^{\prime}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$\scriptstyle{F_{(-)^{\ast}}}$$\scriptstyle{\operatorname{BC}_{2}}$$\scriptstyle{\xi^{2}}$$\scriptstyle{\xi^{\prime}}$
(4.8) $(g_{1}^{\prime})_{\ast}((t^{\prime})^{\ast}(f_{2}^{\prime})^{\ast}(t_{1})_{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}((t^{\prime})^{\ast}t_{\ast}^{\prime}(g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g^{\prime}_{1})_{\ast}((g_{2}^{\prime})^{\ast}t_{1}^{\ast}(t_{1})_{\ast}E\otimes(t^{\prime})^{\ast}K)$$(g_{1}^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K)$$\scriptstyle{\operatorname{BC}_{1}}$$\scriptstyle{F_{(-)^{\ast}}}$$\scriptstyle{\xi^{\prime}}$$\scriptstyle{\xi^{1}}$
Assume that we have shown Diagrams˜4.3, 4.4, 4.5, 4.6, 4.7 and 4.8 are commutative. Then Diagram˜4.2 (in the case of $E$) is the pasting of these diagrams, i.e. gives the desired morphism
$t_{2}^{\ast}(f_{1}^{\prime})_{\ast}((f_{2}^{\prime})_{\ast}(t_{1})_{\ast}E\otimes K)$$(g_{1}^{\prime})_{\ast}((g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K).$}}
So, let us explain why Diagrams˜4.3, 4.4, 4.5, 4.6, 4.7 and 4.8 are commutative. As for Diagram˜4.3, one may use naturality of $\operatorname{BC}_{2}$, whereas Diagram˜4.4 is due to the naturality of $\operatorname{BC}_{2}$. Also, for Diagram˜4.5, one can use the naturality of $M_{(-)^{\ast}}$.
Next, we explain Diagram˜4.6. Observe that it is the functor $(g_{1}^{\prime})_{\ast}$ applied to the following diagram evaluated at $((g_{2}^{\prime})^{\ast}E,K)$:
$(t^{\prime})^{\ast}(t_{\ast}^{\prime}(-)\otimes-)$$(t^{\prime})^{\ast}t_{\ast}^{\prime}(-\otimes(t^{\prime})^{\ast}(-))$$(t^{\prime})^{\ast}t_{\ast}^{\prime}(-)\otimes(t^{\prime})^{\ast}(-)$$(-)\otimes(t^{\prime})^{\ast}(-)$$\scriptstyle{M_{(-)^{\ast}}}$$\scriptstyle{\xi^{\prime}}$$\scriptstyle{\xi^{\prime}}$
which commutes by definition of the projection formula, see e.g. proof of [GW23, Proposition 22.81].
Now, for Diagram˜4.7. Note that it is the following diagram evaluated at $(g_{2}^{\prime})^{\ast}E\otimes(t^{\prime})^{\ast}K$:
[…]
which commutes by definition of the base change morphism. See [GW23, Definition and Remark 21.129] or [LIP09, (3.7.2)].
Lastly, we check Diagram˜4.8. However, it is the functor $(g_{1}^{\prime})_{\ast}(-\otimes(t^{\prime})^{\ast}K)$ applied to the following diagram evaluated at $E$:
[…]
which commutes by definition of the base change morphism. ∎
Remark 4.6.
Assume the hypothesis of Lemma˜4.5. Denote by $\zeta^{i}$ for the unit of $t_{i}^{\ast}\dashv(t_{i})_{\ast}$ if $i\in\{1,2\}$. From Lemmas˜4.1 and 4.5, we have that the following diagram is commutative:
(4.9) $\Phi_{K}$$(t_{2})_{\ast}t_{2}^{\ast}\Phi_{K}$$\Phi_{K}(t_{1})_{\ast}t_{1}^{\ast}$$(t_{2})_{\ast}\Phi_{(t^{\prime})^{\ast}K}t_{1}^{\ast}.$$\scriptstyle{\zeta^{2}_{\Phi_{K}}}$$\scriptstyle{\Phi_{K}(\zeta^{1})}$$\scriptstyle{(t_{2})_{\ast}(\beta^{K})}$$\scriptstyle{\alpha^{K}_{t_{1}^{\ast}}}$
@@
Consider ˜3.1 where $t$ is affine. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Denote by $K^{\prime}$ for the kernel of the integral transform obtained in Proposition˜3.8 which is right adjoint to $\Phi_{K}$ on $D_{\operatorname{qc}}$. Then, after possibly replacing $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ by an isomorphic functor, we have a morphism of adjoint pairs:
$D_{\operatorname{qc}}(Y_{1})$$D_{\operatorname{qc}}(Y_{1}\times_{S}T)$$D_{\operatorname{qc}}(Y_{2})$$D_{\operatorname{qc}}(Y_{2}\times_{S}T)$$\scriptstyle{\mathbf{L}t_{1}^{\ast}}$$\scriptstyle{\Phi_{K}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}}$$\scriptstyle{\Phi_{K^{\prime}}}$$\scriptstyle{\mathbf{L}t_{2}^{\ast}}$$\scriptstyle{\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}}\scriptstyle{\dashv}\scriptstyle{\dashv}$
with left and right comparison transformations given respectively by the isomorphisms
$(\beta^{K})^{-1}\colon\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}\to\mathbf{L}t_{2}^{\ast}\circ\Phi_{K},$
$\beta^{K^{\prime}}\colon\mathbf{L}t_{1}^{\ast}\circ\Phi_{K^{\prime}}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}\circ\mathbf{L}t_{2}^{\ast}.$